Eugene earns a monthly salary of R10 000. He gets a travel allowance of R2000. The employer pays Eugene's medical aid of R500. R400 per month is deducted from his salary for a pension fund. In February he earned commission of R5000. (Eugene is the only member on the medical aid fund. His company contributes R500 while he makes no contribution towards the fund).


Calculate Eugene's Maximum monthly allowable pension fund deduction.


1
Expert's answer
2022-03-24T16:11:29-0400

"max \: deduction= 27.5%\cdot taxable income"% x taxable income

"=\frac{27.5}{100} \times (10000+2000+500+5000)"

=0.275"\times17500"

=4812.50